Mars Society launches Kickstarter to fund a VR platform for the MDRS

The Mars Society launched today a Kickstarter campaign to help raise $27,500 for a new open-source virtual reality platform called MarsVR, which will be used for serious research to support the goal of sending humans to the Red Planet.

The MarsVR program will be a unique multi-phase effort designed to pioneer the emerging field of CrowdExploration, which we define as the partnership between the first astronauts on Mars and VR experts and enthusiasts back on Earth. The Mars Society aims to develop a special VR platform to assist with the initial human exploration of Martian landing sites.

Phase 1 of the MarsVR program will focus on designing training simulations for the Mars Society’s Mars Desert Research Station in Utah, providing direct assistance in preparing MDRS crew members for their analog research and testing. Since every moment at MDRS is valuable, just like actual astronauts working in space, the Mars Society wants its crew members to make the most of their time “on Mars.” Key elements of the platform will also be open-sourced so that the general public can freely make use of it to experience human Mars exploration.

As part of MarsVR, the Mars Society intends to build a high-resolution simulation of the entire MDRS habitat, both inside and out. To complete the experience, Mars Society staff will scan a one-square mile capture of the Mars-like terrain around MDRS using the latest photo-grammetry techniques. Funds raised beyond the $27,500 campaign goal will be used to expand the VR platform beyond the initial one-square mile capture to allow for a broader Mars experience for both crew and members of the public.

For full details about our MarsVR program and how to contribute to the Kickstarter campaign, please visit:

Sol 8 Summary – April 30th

Crew 193 Sol Summary Report – 30 APR 2018

Sol 8

Summary Title: “It was a good Sol” – Ice Cube

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Nominal

Sol Activity Summary: The day started off later than normal as the first EVA was not scheduled to leave until 0800. EVA 14 was able to explore the White Moon area on Mars! While EVA 14 was occurring, the two crew members in the hab took care of the plants in the green Hab and cleaned the shared living area in the main Hab. One crew member credits the growth of the plants to the time she spends talking to them. Maybe Mark Watney should have tried that!

The crew lunched together (thanks again to our commander for cooking), had a short meeting to continue team development and then it was time to get ready for EVA 15. EVA 15 was cut short at 1800 due to approaching rain clouds. The EVA was in the Candor Chasma area and it was decided that if it rained it would be a very difficult area to exit and return to the Hab.

Look Ahead Plan: EVA 16 will be exploring Lith Canyon and the Dinosaur Quarry. EVA 17 will explore the area around Zubrin’s Head and Southwest. We will continue to work on learning ASL for 5 minutes each day as well as sharing stories about our lives so that the crew gets to know each other better. Our daily team meetings have been working well and will be continued through the remainder of the mission.

Anomalies in work:

1. None to report today

Weather: Partly cloudy in AM, cloudy and gusty in PM with rain in the distance.

Crew Physical Status: Minor physical ailments (blister, etc.) but otherwise nominal!


EVA 14 – Explored the moon area. Continued the GPS study and improved on their ATV skills by taking their most challenging road off of Cowdung Road yet.

EVA 15 – Successfully explored the Candor Chasma area and obtained some spectacular photographs. EVA was cut short early due to the possibility of incoming rain making it difficult to exit the Candor Chasma area.

**Please note that the EVA numbers submitted in the EVA report are not correct. It should have been EVA 14 and 15, not 15 and 16**

Reports filed:


Science Report

Astronomy Report

Operations Report

EVA Report

Reports to be filed:

Commander’s Report

Photo & Photo of the Day Report

Next Day EVA Request

Support Requested:

1. Hot Tub? (No requests at this time)

EVA 15 & 16 Summary Reports – April 30th

Sol 8 Crew 193 EVA 15 & 16 Summary
Earth-Date 30APR2018

EVA 15

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold 003 (CE Attig) / Gold005 (CGA Shear)

EVA Objectives – Summary
1. Explore The Moon
2. Repeated GPS recording experiment.
3. Continue to explore the crew biometrics study while on EVA.

Was somewhat windy and chilly in morning. Took a wrong turn (we thougth it was the ATV road), quickly realized and managed to turn around and return to road 1575. No major issues

ATVs/Rovers Used
ATV 300
Honda ATV
ATV #350.1

Safety Measures & Preparation
Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
Adequate hydration – No
Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA 16

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold 004 (HSO Pandya) /

EVA Objectives – Summary
Explore Candor Chasma, try to find marker left by previous expedition.
Continue to explore the crew biometrics study while on EVA
Continued search for emergency shelters

Extreme winds.

ATVs/Rovers Used
Honda ATV
ATV #350.3

Safety Measures & Preparation
Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
Adequate hydration – No
Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

GreenHab Report – April 30th

GreenHab Report

Date: 30Apr2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: not used today

Average temperatures: 24.3 ℃

Low: 15.5℃

High: 50.2℃

Hours of supplemental light: None

Changes to crops: Yellowing resolving, plants seem to be in slightly better health.

Daily water usage for crops: 7.25 Gallons

Time(s) of watering for crops: Twice: In the morning at 0800 and at 1845.

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: Yellowed plants are are turning green again after extra watering. It may or may not help that we are talking to them gently but lovingly.

Support/supplies needed: None

Operations Report – April 30th

Crew 193 Operations Report 30Apr2018

SOL: 8

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now). At 4:45 PM today, the SoC was 73%, which is lower than it should be, given how sunny it is today; there are some storm clouds to the southwest, which could be blocking the sun, we need to keep an eye on SoC.
I will repair charging plugs once we have a soldering iron.

Generator (hours run): 10:45 pm to 8:10 am.

Solar— SOC 59 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134370.9

Toilet tank emptied: no
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 3

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding.

Summary of RAM operations: A charging connector was moved from RAM for space suit repair (once soldering iron arrives)

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: We would like some black straps used for connecting two piece suit neck rings to backpacks. I did not see any in the RAM.

Science Report – April 30th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters. The second EVA found a location in Candor Chasma that had been previously scouted by the other EVA team last week.

2. Sunspot and Solar Flare Monitoring: The crew astronomer saw one filament on the sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Commander’s Report – April 30th

Commander, PHEnOM Gold Crew, MDRS #193

Mission Sol 8 (Earth Day April 30th, 2018)

Exactly a week ago we arrived at Mars and embarked on this mission, excited, curious, and eager to learn, explore, research and contribute to mankind’s endeavor to colonize Mars. It has been a rewarding week what with all our research progressing well, exploration and scouting for emergency shelters giving positive results and barring a few minor incidents, the crew staying fit and fine.

The crew has adapted well to the harsh Martian conditions and is functioning great as one unit. I am impressed at how each member is being creative and innovative in coming up with solutions to problems and willing to take on new projects to help in every way to make this mission a success.

Mars is challenging ; its comforting to see the crew maintain a stable emotional quotient and in great psychological health even after a week here! Spirits remain high and the crew looks forward to completing their mission goals and more.

We have been working hard and playing hard. Games, movies, a lot of sharing stories and learning ASL have been our fun pastimes. Our crew is a great example of how people from different backgrounds, age groups and professions can come together and work as a team to make a mission successful.

Every sol on Mars has been an adventure to behold,
For this outstanding crew, call us 193 or Gold!
With all the work we do here, we’ll leave our mark on Mars,
As we continue to further mans mission, to travel to the stars !
– Ur Poet Commander 😉

Journalist Report – April 29th

Journalist Report – Earth-Date 29APR2018 / Mars Sol 7

Author: Eric Shear

This weekend, we took a much-needed rest day on Sol 6. That doesn’t mean we lounged around all day – we made outreach videos with a combination of spoken and signed language.

Today (Sol 7) began the first of our switched EVAs. Team Canada had previously done morning EVAs, and is now doing EVAs in the afternoon, and vice versa for the Americans.

On Sol 6, our geologist/astronomer finally caught a prominence on the Sun, much to the excitement of the astronomers at Mission Support. Then he did it again on Sol 7, getting three solar features despite heavy cloud cover. We are glad to see him sharpen his already considerable telescope and image processing skills.

The morning of Sol 7, Team America went on EVA to Zubrin’s Head, White Rock Canyon, and Kissing Camel. Team Canada explored Toothy Ridge and Copernicus Highway. During one of these EVAs, a crewmember injured her knee by falling down a hill.

The afternoon was spent recuperating, working on projects and furthering our research. A quiet evening of eating good food and low-key recreation is anticipated. Our XO is anxious for a rematch in cards…

Copyright © The Mars Society. All rights reserved. | Main Site